# Nonlinear Magnetically Charged Black Holes with Phantom Global Monopoles: Thermodynamics, Geodesics, Quasinormal Modes, and Grey-Body Factors

### v2

@@ 4 Timelike Geodesics
[…]
@@
Eq. (40) describes a system analogous to the motion of a classical particle with energy $E^{2}$ in an effective one-dimensional potential $V_{\mathrm{eff}}(r)$. The nature of this potential determines key features of the geodesic motion, such as stable circular orbits, escape conditions, and critical impact parameters for light bending. The influence of the phantom field, nonlinear electrodynamics, and the global monopole manifests in modifications to this potential, altering the behavior of both massive and massless test particles.
@@
[…]
@@
[…]
@@
$0\left.11Q^{3}r^{4}(1-8\pi\eta^{2}\xi)-8Q^{6}r(1-8\pi\eta^{2}\xi)+r^{6}(r-6-8\pi\eta^{2}\xi r)\right|_{r=r_{\mathrm{ISCO}}}=$ 0 . (61)
@@ 5 Quasinormal Modes
Quasinormal modes characterize the response of a black hole to external perturbations and provide crucial insights into its stability and observational signatures in gravitational wave astronomy [69, 70]. The spectrum of these modes depends on the black hole parameters and surrounding matter fields, making them valuable probes of deviations from general relativity.
@@
Additionally, NED introduces further modifications to the QNM spectrum by altering the effective potential governing perturbations. These effects have been investigated in contexts such as Einstein-Gauss-Bonnet gravity [80], AdS black holes with Coulomb-like corrections [81], and charged black holes with nonlinear electrodynamics-induced modifications [82, 83]. Studies have also examined how these effects influence black hole shadows and grey-body factors [84, 85], while additional work on Euler-Heisenberg black holes suggests notable shifts in the QNM spectrum [86].
In this section, we analyze the quasinormal modes of a nonlinear magnetic-charged black hole with a phantom global monopole. Using the sixth-order WKB approximation, we compute the mode spectrum and compare our results with the Pöschl-Teller potential approach to assess accuracy. Our goal is to determine how the interplay between the phantom global monopole and nonlinear electrodynamics influences the black hole’s dynamical response and stability under scalar perturbations.
@@
The behavior of the quasinormal modes is determined by imposing physically motivated boundary conditions on Eq. (69). These conditions ensure purely ingoing waves at the black hole horizon and purely outgoing waves at spatial infinity, leading to:
@@
The effective potential $\mathcal{V}_{\mathrm{eff}}(x)$ determines the QNM spectrum, as it represents the potential barrier through which perturbations must tunnel before being absorbed by the black hole. Its shape and peak height influence the stability and damping rates of the modes. In the following, we analyze the numerical methods used to compute the QNM spectrum and discuss the effects of the phantom global monopole and nonlinear electrodynamics on the stability of the black hole.
@@ 5.1 Computation of QNMs using the WKB Approximation
To compute the quasinormal frequencies, we employ the WKB approximation, a widely used semi-analytical method in black hole physics [89, 90, 87, 88, 94, 95, 96, 86, 97, 98, 93, 91, 92]. Originally introduced at first order in [99], the method was later extended to higher orders, including second and third order in [100] and up to sixth order in [101]. While higher-order corrections enhance precision, their effectiveness depends on the potential shape and parameter regime, as discussed in the reviews [70, 102].
@@
Since Eq. (71) involves nontrivial dependencies on physical parameters, we numerically compute the QNM frequencies for different parameter values. However, as the WKB approximation is most reliable when $\ell>n$, we restrict our analysis to scalar field perturbations satisfying this condition, corresponding to the low-lying quasinormal modes. Selected results for these modes are presented in Table 2
@@
The numerical results given in Tables 2 and 3 underlines the several key results:
@@
Figure 9 illustrates an example of QNMs computed by using the sixth-oreder WKB formula for scalar perturbations with $\eta=0.12$, $\xi=-1$ and $Q=0.3$.
@@
As evident from the figure, the accuracy of the WKB formula decreases for scalar perturbations when $n\geq l$, whereas the results converge more reliably for cases where $l>n$.
@@ 6 Grey-body factors
Grey-body factors and quasinormal modes can be computed using the higher-order WKB. The formulas were established for spherically symmetric and asymptotically flat black holes by applying the well-known WKB method to calculate grey-body factors,
@@
Here $\omega_{0}$ is the the fundamental mode, and $\omega_{1}$ is the first overton. Figure 11 illustrates the grey-body factors as a function of frequency $\Omega$ for a nonlinear magnetic-charged black hole with both ordinary ($\xi=1$) and phantom ($\xi=-1$) global monopoles.
@@
Figure 11: Grey-body factors for $\ell=3$ and $Q=0.1$
@@
Figure 12: Grey-body factors for $\ell=3$ and $\eta=0.1$
@@
Figure 13: Grey-body factors for $\ell=3$, $\eta=0.1$ and $Q=0.2$
As seen in the figure, increasing $Q$ leads to a reduction in the transmission coefficient, meaning that a stronger magnetic charge suppresses the escape of radiation. This behavior is expected since a larger $Q$ enhances the curvature effects and modifies the effective potential, making it more difficult for radiation to overcome the potential barrier. Notably, this effect is consistent across both ordinary and phantom monopole cases, reinforcing the role of nonlinear electrodynamics in shaping black hole radiation characteristics.
@@ 7 Conclusion
In this work, we investigated the properties of a nonlinear magnetic-charged black hole in the presence of a phantom global monopole. By incorporating nonlinear electrodynamics (NED) and exotic scalar fields, we derived an exact black hole solution and analyzed its thermodynamic and dynamical properties. Our findings reveal significant deviations from standard general relativity (GR), providing potential observational signatures for modified gravity theories.
@@
We examined the black hole’s thermodynamic properties, including its Hawking temperature, entropy, and heat capacity. The presence of the phantom global monopole was shown to modify the thermodynamic stability, leading to novel phase structures. Notably, the Hawking temperature exhibits a dependence on the monopole parameter, influencing the evaporation process of the black hole. The heat capacity analysis indicated the presence of a second-order phase transition, suggesting regions of stability and instability in the black hole’s thermodynamic phase space.
@@
We also analyzed the geodesic structure of test particles, exploring their motion in the modified black hole background. Furthermore, we studied the quasinormal modes (QNMs) of the black hole, employing both the sixth-order WKB approximation and the Pöschl-Teller potential method. The QNM spectrum provided insights into the stability of the system under perturbations. Our results showed that the presence of nonlinear electrodynamics and the phantom global monopole modifies the real and imaginary parts of the QNM frequencies, affecting the damping times of perturbations.
@@
A key addition to our study was the analysis of grey-body factors, which describe the probability of radiation escaping the black hole’s gravitational potential. We found that the presence of the phantom global monopole reduces the transmission coefficient, leading to a suppression of escaping radiation. This effect was further influenced by the symmetry-breaking parameter and the magnetic charge, reinforcing the role of nonlinear electrodynamics in shaping black hole radiation properties.
@@
Overall, our findings highlight the intricate interplay between nonlinear electrodynamics, phantom fields, and global monopoles in black hole physics. The modifications introduced by these exotic fields have direct implications for astrophysical observations, including black hole shadows, gravitational waves, and Hawking radiation. Future work could extend this analysis by exploring rotating solutions, higher-dimensional generalizations, or the impact of additional modified gravity corrections.
@@ Acknowledgments
This work is supported by the Ministry of Higher Education and Scientific Research, Algeria under the code: B00L02UN040120230003. B. C. L. is grateful to Excellence Project PřF UHK 2205/2025-2026 for the financial support.

### v3

@@ 4 Timelike Geodesics
A notable property of geodesic motion in spherically symmetric spacetimes is that if a particle initially moves within the equatorial plane ($\theta=\pi/2$ at an initial time $t_{i}$), it remains confined to this plane throughout its motion. This follows directly from the equation for $\ddot{\theta}$, which vanishes under these conditions. As a result, the motion can be analyzed without loss of generality by restricting to the equatorial plane, leading to the simplified geodesic equations:
@@
Eq. (45) describes a system analogous to the motion of a classical particle with energy $E^{2}$ in an effective one-dimensional potential $V_{\mathrm{eff}}(r)$. The nature of this potential determines key features of the geodesic motion, such as stable circular orbits, escape conditions, and critical impact parameters for light bending. The influence of the phantom field, NLE, and the global monopole manifests in modifications to this potential, altering the behavior of both massive and massless test particles.
@@
$\frac{\partial^{2}V_{\mathrm{eff}}(r)}{\partial r^{2}}>0\quad$ $\Rightarrow$ $\quad\text{Unstable},$ (55)
$\frac{\partial^{2}V_{\mathrm{eff}}(r)}{\partial r^{2}}<0\quad$ $\Rightarrow$ $\quad\text{Stable}.$ (56)
@@
$\frac{\partial^{2}V_{\mathrm{eff}}(r)}{\partial r^{2}}>0\quad$ $\Rightarrow$ $\quad\text{Unstable},$ (62)
$\frac{\partial^{2}V_{\mathrm{eff}}(r)}{\partial r^{2}}<0\quad$ $\Rightarrow$ $\quad\text{Stable}.$ (63)
@@
$\left.11Q^{3}r^{4}(1-8\pi\eta^{2}\xi)-8Q^{6}r(1-8\pi\eta^{2}\xi)+r^{6}(r-6-8\pi\eta^{2}\xi r)\right|_{r=r_{\mathrm{ISCO}}}=0.$ (66)
@@ 5 Quasinormal Modes
QNMs characterize the response of a black hole to external perturbations and provide crucial insights into its stability and observational signatures in gravitational wave astronomy [72, 73]. The spectrum of these modes depends on the black hole parameters and surrounding matter fields, making them valuable probes of deviations from general relativity.
@@
Additionally, NLE introduces further modifications to the QNM spectrum by altering the effective potential governing perturbations. These effects have been investigated in contexts such as Einstein-Gauss-Bonnet gravity [83], AdS black holes with Coulomb-like corrections [84], and charged black holes with NLE-induced modifications [85, 86]. Studies have also examined how these effects influence black hole shadows and grey-body factors [87, 88], while additional work on Euler-Heisenberg black holes suggests notable shifts in the QNM spectrum [89].
In this section, we analyze the QNMs of a nonlinear magnetic-charged black hole with a phantom global monopole. Using the sixth-order WKB approximation, we compute the mode spectrum and compare our results with the Pöschl-Teller potential approach to assess accuracy. Our goal is to determine how the interplay between the phantom global monopole and NLE influences the black hole’s dynamical response and stability under scalar perturbations.
@@
The behavior of the QNMs is determined by imposing physically motivated boundary conditions on Eq. (74). These conditions ensure purely ingoing waves at the black hole horizon and purely outgoing waves at spatial infinity, leading to:
@@
The effective potential $\mathcal{V}_{\mathrm{eff}}(x)$ determines the QNM spectrum, as it represents the potential barrier through which perturbations must tunnel before being absorbed by the black hole. Its shape and peak height influence the stability and damping rates of the modes. In the following, we analyze the numerical methods used to compute the QNM spectrum and discuss the effects of the phantom global monopole and NLE on the stability of the black hole.
@@ 5.1 Computation of QNMs using the WKB Approximation
To compute the quasinormal frequencies, we employ the WKB approximation, a widely used semi-analytical method in black hole physics [92, 93, 90, 91, 97, 98, 99, 89, 100, 101, 96, 94, 95, 102, 103]. Originally introduced at first order in [104], the method was later extended to higher orders, including second and third order in [105] and up to sixth order in [106]. While higher-order corrections enhance precision, their effectiveness depends on the potential shape and parameter regime, as discussed in the reviews [73, 107].
@@
Since Eq. (76) involves nontrivial dependencies on physical parameters, we numerically compute the QNM frequencies for different parameter values. However, as the WKB approximation is most reliable when $\ell>n$, we restrict our analysis to scalar field perturbations satisfying this condition, corresponding to the low-lying QNMs. Selected results for these modes are presented in Table 3,
@@
The numerical results given in Tables 3 and 4 underlines the several key results:
@@
Figure 9 illustrates an example of QNMs computed by using the sixth-order WKB formula for scalar perturbations with $\eta=0.12$, $\xi=-1$ and $Q=0.3$.
@@
As evident from the figure, the accuracy of the WKB formula decreases for scalar perturbations when $n\geq{\ell}$, whereas the results converge more reliably for cases where ${\ell}>n$.
@@ 6 Grey-body factors
Grey-body factors and QNMs can be computed using the higher-order WKB. The formulas were established for spherically symmetric and asymptotically flat black holes by applying the well-known WKB method to calculate grey-body factors,
@@
Here, $\omega_{0}$ is the the fundamental mode and $\omega_{1}$ is the first overton. Figure 11 illustrates the grey-body factors as a function of frequency $\Omega$ for a nonlinear magnetic-charged black hole with both ordinary ($\xi=1$) and phantom ($\xi=-1$) global monopoles.
@@
Figure 11: Grey-body factors for $\ell=3$ and $Q=0.1$.
@@
Figure 12: Grey-body factors for $\ell=3$ and $\eta=0.1$.
@@
Figure 13: Grey-body factors for $\ell=3$, $\eta=0.1$ and $Q=0.2$.
As seen in the figure, increasing $Q$ leads to a reduction in the transmission coefficient, meaning that a stronger magnetic charge suppresses the escape of radiation. This behavior is expected since a larger $Q$ enhances the curvature effects and modifies the effective potential, making it more difficult for radiation to overcome the potential barrier. Notably, this effect is consistent across both ordinary and phantom monopole cases, reinforcing the role of NLE in shaping black hole radiation characteristics.
@@ 7 Conclusion
In this work, we investigated the properties of a nonlinear magnetic-charged black hole in the presence of a phantom global monopole. By incorporating nonlinear electrodynamics (NLE) and exotic scalar fields, we derived an exact black hole solution and analyzed its thermodynamic and dynamical properties. Our findings reveal significant deviations from standard general relativity (GR), providing potential observational signatures for modified gravity theories.
@@
We examined the black hole’s thermodynamic properties, including its Hawking temperature, entropy, and heat capacity. The presence of the phantom global monopole was shown to modify the thermodynamic stability, leading to novel phase structures. Notably, the Hawking temperature exhibits a dependence on the monopole parameter, influencing the evaporation process of the black hole. The heat capacity analysis indicated the presence of a second-order phase transition, suggesting regions of stability and instability in the black hole’s thermodynamic phase space. Although the influence of the phantom monopole may appear as an overall shift in the temperature and heat capacity curves, this shift encodes significant physical effects: the conical excess introduced by the phantom field alters the asymptotic geometry, reduces the black hole temperature at fixed horizon radius, and shifts the location of phase transition points. These changes reflect a genuine modification of the black hole’s thermodynamic behavior due to the topological deformation introduced by the scalar field.
@@
We also analyzed the geodesic structure of test particles, exploring their motion in the modified black hole background. The analysis reveals that both the phantom global monopole and the nonlinear magnetic charge contribute significantly to the structure and stability of circular orbits. Notably, the presence of the monopole tends to shift the innermost stable circular orbit (ISCO) outward and enhance orbital instability, while the NLE acts to stabilize orbits at smaller radii. The Lyapunov exponent, which measures the instability timescale of orbits, increases with the symmetry-breaking parameter $\eta$, mirroring the faster damping rates observed in the quasinormal mode spectrum. These findings not only reinforce the consistency between the dynamical and perturbative analyses, but also suggest possible observational signatures in accretion disk dynamics and gravitational wave signals near such exotic black holes. Furthermore, we studied the quasinormal modes (QNMs) of the black hole, employing both the sixth-order WKB approximation and the Pöschl-Teller potential method. The QNM spectrum provided insights into the stability of the system under perturbations. Our results showed that the presence of NLE and the phantom global monopole modifies the real and imaginary parts of the QNM frequencies, affecting the damping times of perturbations.
@@
A key addition to our study was the analysis of grey-body factors, which describe the probability of radiation escaping the black hole’s gravitational potential. We found that the presence of the phantom global monopole reduces the transmission coefficient, leading to a suppression of escaping radiation. This effect was further influenced by the symmetry-breaking parameter and the magnetic charge, reinforcing the role of NLE in shaping black hole radiation properties.
@@
Overall, our findings highlight the intricate interplay between NLE, phantom fields, and global monopoles in black hole physics. The modifications introduced by these exotic fields have direct implications for astrophysical observations, including the shadows of black holes, gravitational waves, and Hawking radiation. Future work could extend this analysis by exploring rotating solutions, higher-dimensional generalizations, or the impact of additional modified gravity corrections.
@@ Acknowledgments
The authors thank the anonymous reviewers for their constructive comments and valuable suggestions, which greatly improved the quality of this work. B. C. L. is grateful to Excellence Project FoS UHK 2205/2025-2026 for the financial support.